# Bayesian Optimization for Controlled Image Editing via LLMs

### v2

@@ Abstract
In the rapidly evolving field of image generation, achieving precise control over generated content and maintaining semantic consistency remain significant limitations, particularly concerning grounding techniques and the necessity for model fine-tuning. To address these challenges, we propose BayesGenie, an off-the-shelf approach that integrates Large Language Models (LLMs) with Bayesian Optimization to facilitate precise and user-friendly image editing. Our method enables users to modify images through natural language descriptions without manual area marking, while preserving the original image’s semantic integrity. Unlike existing techniques that require extensive pre-training or fine-tuning, our approach demonstrates remarkable adaptability across various LLMs through its model-agnostic design. BayesGenie employs an adapted Bayesian optimization strategy to automatically refine the inference process parameters, achieving high-precision image editing with minimal user intervention. Through extensive experiments across diverse scenarios, we demonstrate that our framework significantly outperforms existing methods in both editing accuracy and semantic preservation, as validated using different LLMs including Claude3 and GPT-4.
@@
Chengkun Cai1∗,  Haoliang Liu2∗,  Xu Zhao1∗,  Zhongyu Jiang3,  Tianfang Zhang4, Zongkai Wu5,  Jenq-Neng Hwang3,  Serge Belongie6,  Lei Li3,6† thanks: * Equal contribution. thanks: $\dagger$ Corresponding author: lilei@di.ku.dk 1University of Edinburgh  2University of Manchester  4Tsinghua University 5FancyTech  3University of Washington  6University of Copenhagen
@@ 1 Introduction
However, despite these advancements, existing methods face several critical challenges. First, most state-of-the-art local editing methods heavily rely on mask priors to constrain the editing regions—either through manual input or derived from attention map analysis and semantic segmentation—making them less accessible for non-expert users. Additionally, methods like cross-attention control and diffusion models often encounter challenges in fine-tuning model parameters to align with user requirements, resulting in a disparity between desired and actual outputs. These issues are particularly pronounced in applications that demand detailed modifications based on user instructions. With the LLMs-Driven adaptation, recent various applications Cai et al. (2024b, a); Li et al. (2024a); Shi et al. (2025); Liu et al. (2024); Shi et al. (2024) have advanced developed.
@@
  • We propose BayesGenie, a novel image editing framework that enables precise localized editing without manual region annotations or mask creation. Our model-agnostic approach leverages LLMs for region understanding and semantic interpretation, making it possible to perform accurate local edits based purely on natural language descriptions, while ensuring high-precision image editing.
@@ 1.1 Related Work
However, a common limitation persists across most state-of-the-art approaches: they typically rely on some form of mask guidance to achieve precise local editing. Whether through manual mask annotations Hertz et al. (2022), attention map analysis Li et al. (2024b), or semantic segmentation, the dependence on mask priors creates a barrier for non-expert users and limits the flexibility of these systems. BayesGenie addresses this limitation by enhancing the diffusion model with automated parameter optimization, enabling it to follow instructions more accurately while preserving the high quality of the generated images without requiring any form of mask guidance.
@@ Appendix B Scoring Evaluation Prompt
Appendix C Ethics and Participant Consent

### v3

@@ Abstract
In the rapidly evolving field of image generation, achieving precise control over generated content and maintaining semantic consistency remains a crucial limitation, particularly concerning grounding techniques and the necessity for model fine-tuning. To address these challenges, we propose BayesGenie, an off-the-shelf approach that integrates Large Language Models (LLMs) with Bayesian Optimization to facilitate precise and user-friendly image editing. Our method enables users to modify images through natural language descriptions without manual area marking, while preserving the original image’s semantic integrity. Unlike existing techniques that require extensive pre-training or fine-tuning, our approach demonstrates remarkable adaptability across various LLMs through its model-agnostic design. BayesGenie employs an adapted Bayesian optimization strategy to automatically refine the inference process parameters, achieving high-precision image editing with minimal user intervention. Through extensive experiments across diverse scenarios, we demonstrate that our framework outperforms existing methods in both editing accuracy and semantic preservation, as validated using different LLMs including Claude3 and GPT-4.
@@
Chengkun Cai1thanks: Equal contribution.  Haoliang Liu211footnotemark: 1  Xu Zhao [^1] footnotemark: 1  Zhongyu Jiang3  Tianfang Zhang4 Zongkai Wu5thanks: Project Leader.  John Lee1  Jenq-Neng Hwang3  Lei Li3,6thanks: Corresponding author: lilei@di.ku.dk 1University of Edinburgh  2University of Manchester  3University of Washington 4Tsinghua University  5Skai Intelligence  6University of Copenhagen
@@ 1 Introduction
However, despite these advancements, existing methods face several critical challenges. First, most state-of-the-art local editing methods heavily rely on mask priors to constrain the editing regions—either through manual input or derived from attention map analysis and semantic segmentation—making them less accessible for non-expert users. Additionally, methods like cross-attention control and diffusion models often encounter challenges in fine-tuning model parameters to align with user requirements, resulting in a disparity between desired and actual outputs. These issues are particularly pronounced in applications that demand detailed modifications based on user instructions. With the LLMs-Driven adaptation, recent various applications Cai et al. (2024a); Li et al. (2024a); Shi et al. (2025); Liu et al. (2024); Shi et al. (2024); Cai et al. (2024b) have advanced developed.
@@
  • We propose BayesGenie, a novel image editing framework that enables precise localized modifications without the need for manual region annotations or mask generation. The framework adopts a model-agnostic design that leverages large language models (LLMs) for spatial reasoning and semantic interpretation, allowing for accurate and context-aware image editing based solely on natural language instructions. Distinctively, BayesGenie is both training-free and mask-free, which endows it with strong generalization capabilities and high adaptability. This design choice allows BayesGenie to be seamlessly integrated with a wide range of existing image editing backbones—regardless of whether the underlying models are mask-based, fine-tuned, or operate in a zero-shot manner.
@@ 1.1 Related Work
However, a common limitation persists across most state-of-the-art approaches: they typically rely on some form of mask guidance to achieve precise local editing. Whether through manual mask annotations Hertz et al. (2022), attention map analysis Li et al. (2024b), or semantic segmentation, the dependence on mask priors creates a barrier for non-expert users and limits the flexibility of these systems. Recent works have begun to overcome this challenge through fully mask-free editing mechanisms. For instance, strong mask-free image editing baselines have emerged, such as MagicBrush Zhang et al. (2023) and UltraEdit Zhao et al. (2024), which enable high-quality, instruction-driven editing without relying on explicit masks.
@@ 5 Conclusion
6 Ackonwledgement
This work was supported in part by the Pioneer Centre for AI, DNRF grant number P1.
@@ Appendix B Scoring Evaluation Prompt
Appendix C Comparison with State-of-the-Art Models
To further validate the effectiveness and generalizability of our proposed BayesGenie framework, we conducted additional experiments comparing it with several recent state-of-the-art image editing models, including MagicBrush Zhang et al. (2023) and UltraEdit Zhao et al. (2024). The evaluation metrics include CLIPScore and GPT-4o-based instruction consistency scores. Results are summarized in Table 2.
Table 2: Performance comparison with state-of-the-art models.
Model CLIPScore GPT-4o Score
InstructPix2Pix 0.2712 23.5
InstructPix2Pix + BayesGenie 0.3180 78.6
MagicBrush 0.3078 53.7
UltraEdit 0.3302 62.6
UltraEdit + BayesGenie 0.3524 85.3
To enable a fair and consistent comparison, we integrated BayesGenie into both the InstructPix2Pix and UltraEdit pipelines. InstructPix2Pix and UltraEdit both support classifier-free guidance (CFG), which BayesGenie optimizes during inference. For UltraEdit, which typically uses spatial masks, we covered the entire image as the editable region to align it with the mask-free nature of BayesGenie.
We evaluated each model on 21 diverse editing tasks (including addition, removal, and modification of objects), with each task repeated five times to report averaged scores. It is worth noting that both MagicBrush and InstructPix2Pix utilize Stable Diffusion 1.5 as their backbone, whereas UltraEdit is based on Stable Diffusion 3, which partially accounts for some of the performance differences observed.
The results highlight two core strengths of the BayesGenie framework:
  1. 1.
    Complementary Nature: BayesGenie significantly improves the performance of existing models such as InstructPix2Pix and UltraEdit, not by competing with them, but by acting as an effective enhancement module.
  2. 2.
    Universal Applicability: BayesGenie demonstrates strong performance across a wide spectrum of architectures, from general-purpose models like InstructPix2Pix to highly specialized systems like UltraEdit. This underscores its versatility and model-agnostic design.
Appendix D Ethics and Participant Consent